Tag: red (page 1 of 31)

Special Message, Final Event Energies Update 6~14~17: The Red Dragon & Solstice Energies ~ Love Has

View Article Here   Read More

The Portal 2017-06-14 08:59:00

Alert downgrade to red

View Article Here   Read More

The Portal 2017-06-11 04:30:00

Alert downgrade to red, BSMOSS final sequence in progress, BSR sequence initialized

View Article Here   Read More

UFO Headline News Friday June 9th, 2017

In case you missed it here is the UFO Headline News for today

The post UFO Headline News Friday June 9th, 2017 appeared first on Inception Radio Network | UFO & Paranormal Talk Radio.

View Article Here   Read More

The Portal 2017-06-02 16:22:00

Isidic security breach deflected, systems red alert cont 504

View Article Here   Read More

The Portal 2017-06-02 16:22:00

Isidic security breach deflected, systems red alert cont 504

View Article Here   Read More

STATE OF MISSION REPORT MAY 2017

Pandora in progress, EELA in progress, BSMOSS sequence in progress. Systems/Isidic security breach deflection in progress, red alert at 504. HVBN  stable. M minimum requirements met.

View Article Here   Read More

STATE OF MISSION REPORT MAY 2017

Pandora in progress, EELA in progress, BSMOSS sequence in progress. Systems/Isidic security breach deflection in progress, red alert at 504. HVBN  stable. M minimum requirements met.

View Article Here   Read More

Gaia Portal ~Assistance Comes As Dark Hu Beings Require ~ May 11 2017

View Article Here   Read More

Gaia Portal ~ Red Flowers Unfold And Illuminate The Hu Masses ~ May 05 2017

View Article Here   Read More

STATE OF MISSION REPORT APRIL 2017

Pandora in progress, EELA in progress. Potentially fatal Systems/Isidic security breach, grid ratio failure, Morph decrease from 150 to 126, deflection attempt in progress. Black alert/red alert at 504. HVBN  stable. M minimum requirements met.

View Article Here   Read More

STATE OF MISSION REPORT APRIL 2017

Pandora in progress, EELA in progress. Potentially fatal Systems/Isidic security breach, grid ratio failure, Morph decrease from 150 to 126, deflection attempt in progress. Black alert/red alert at 504. HVBN  stable. M minimum requirements met.

View Article Here   Read More

The Portal 2017-04-26 17:05:00

Alert downgrade to red at 504

View Article Here   Read More
Older posts




Gaia-Cosmic Disclosure S1E1 LB728x90

Creative Commons License
This work is licensed under a
Creative Commons Attribution 4.0
International License
,
unless otherwise marked.

Terms of Use | Privacy Policy

Member of The Internet Defense League




Up ↑